# Don’t flatten, tokenize! Unlocking the key to SoftMoE’s efficacy in deep RL

### v1

@@ 7 Related Works
Mixture of Experts (MoEs) have demonstrated remarkable success in scaling language and computer vision models to trillions of parameters (Lepikhin et al., 2020; Fedus et al., 2022; Wang et al., 2020; Yang et al., 2019; Abbas and Andreopoulos, 2020; Pavlitskaya et al., 2020). Chung et al. (2024) examined the implicit biases arising from token combinations in SoftMoE and defined the notion of expert specialization in computer vision. Besides the success of MoEs in single task training, similar gains have been observed in transfer and multitask learning (Puigcerver et al., 2020; Chen et al., 2023; Ye and Xu, 2023; Hazimeh et al., 2021). In addition to the works of Obando Ceron* et al. and Willi* et al. already discussed, Farebrother et al. (2024) proposed replacing regression loss with classification loss and showed its benefits on MoEs.
@@
Scaling reinforcement learning (RL) networks presents significant challenges, largely due to issues with training instabilities (Hessel et al., 2018). To address these difficulties, several techniques have been proposed, with a focus on algorithmic improvements or hyperparameter optimization (Farebrother et al., 2022, 2024; Taiga et al., 2022; Schwarzer* et al., 2023). Recently, a few studies have begun to focus on architectural innovations as a means of addressing these challenges. Kumar et al. (2023a) showed that incorporating a learned spatial embedding with the output of a ResNet encoder, prior to flattening it, improves the scalability in offline RL. Obando Ceron et al. (2024) demonstrated that gradual magnitude pruning enables the scaling of value-based agents, resulting in substantial performance gains.
@@
It has been observed that networks trained on non-stationary distributions tend to lose their expressivity and ability to adapt to new data over time, a phenomenon referred to as the loss of plasticity (Kumar et al., 2021; Lyle et al., 2021). While several studies have attempted to explain this phenomenon (Lyle et al., 2023; Ma et al., 2023; Lyle et al., 2024), the precise underlying causes remain uncertain. In a parallel line of research, numerous methods have been proposed to mitigate plasticity loss and improve network trainability. These approaches involve using small batch size for gradient updates (Ceron et al., 2023), periodically resetting some of the network’s parameters to random values (Sokar et al., 2023; Nikishin et al., 2022; Igl et al., 2020; Dohare et al., 2024), slightly modifying the existing weights by shrinking and perturbing them (D’Oro et al., 2022; Schwarzer* et al., 2023), and restricting the large deviation from initial weights (Kumar et al., 2023b; Lewandowski et al., 2024). Recently, Lee et al. have also demonstrated that combining reinitialization with a teacher-student framework can enhance plasticity. Another technique involves preserving useful neurons by perturbing gradients to ensure smaller updates for important neurons while allowing larger updates for less important ones (Elsayed and Mahmood, ). In this work, we explore some of these methods to improve plasticity and utilization of experts.
@@ 8 Conclusions
We began our investigation by evaluating the components that make up a SoftMoE architecture, so as to better understand the strong performance gains reported by Obando Ceron* et al. (2024) when applied to value-based deep RL agents. One of the principal, and rather striking, findings is that tokenizing the encoder outputs is one of the primary drivers of the performance improvements. Since tokenization is mostly necessary for pixel-based environments (where encoders typically consist of convolutional layers), it can explain the absence of performance gains observed by Obando Ceron* et al. (2024) and Willi* et al. (2024) when using MoE architectures in actor-critic algorithms, since these were evaluated in non-pixel-based environments.
More generally, our findings suggests that the common practice of flattening the outputs of deep RL encoders is sub-optimal, as it is likely not maintaining the spatial structure output by the convolutional encoders. Indeed, the results with PerPatch and Shuffled tokenization in Figure 7 demonstrate that maintaining this spatial structure is important for performance. This is further confirmed by our results in Figure 5, where we can observe that not flattening the convolutional outputs can lead to significant improvements. This idea merits further investigation, as the results vary depending on the architecture used and whether features are summed or averaged over.
@@ Acknowledgements
The authors would like to thank Gheorghe Comanici, Jesse Farebrother, Joan Puigcerver, Doina Precup, and the rest of the Google DeepMind team for valuable feedback on this work. We would also like to thank the Python community (Van Rossum and Drake Jr, 1995; Oliphant, 2007) for developing tools that enabled this work, including NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007) and JAX (Bradbury et al., 2018).
@@ B.1 Comparison between token choice routing and expert choice routing
Figure 12 demonstrates that expert choice routing consistently outperforms token choice routing across all numbers of experts. Moreover, this performance gap widens as the number of experts increases, highlighting the advantages of expert choice routing in addressing the limitations of token choice routing.
@@ B.2 Don’t flatten tokenize
In this section, we provide further analysis on the impact of tokenization and their weight combination on models with larger number of experts. To this end, we investigate SoftMoE-8 and compare it against the scaled SoftMoE-1 and ExpertChoice-1 (non-combined tokenization). Figure 13 shows that single scaled expert can reach a performance close to 8 experts. In addition, consistent with our results, combined tokenization adds benefits over the sparse one. Figure 14 shows a comparison between the learning curves of single scaled expert and multiple ones.
Figure 13: Combined tokenization plays a crucial role in performance. SoftMoE-1 with scaled hidden layer achieves a comparable performance to SoftMoE-8.
@@
B.3 More agents
@@ B.4 Expert Utilization
Some experts are redundant
In Section 4.2, we showed that expert are not specialized in learning subset of tokens. We expand this analysis by assessing the redundancy of experts. To this end, we prune two experts of SoftMoE-4 during training. We studied two pruning schemes: pruning the two experts once and gradually prune throughout training. As shown in Figure 16, pruning these experts does not change the performance of the agent.
@@
In investigating current techniques to improve expert utilization, we explore reset either experts weights only or both experts and router weights. As illustrated in Figure 17, both methods do not help in improving utilization, with resetting the router weights leads to worse performance.
Figure 17: Comparison between resetting experts weights only versus experts and router weights in Rainbow (top) and DER (bottom).

### v2

@@ 7 Related Works
Mixture of Experts (MoEs) have demonstrated remarkable success in scaling language and computer vision models to trillions of parameters (Lepikhin et al., 2020; Fedus et al., 2022; Wang et al., 2020; Yang et al., 2019; Abbas & Andreopoulos, 2020; Pavlitskaya et al., 2020). Chung et al. (2024) examined the implicit biases arising from token combinations in SoftMoE and defined the notion of expert specialization in computer vision. Besides the success of MoEs in single task training, similar gains have been observed in transfer and multitask learning (Puigcerver et al., 2020; Chen et al., 2023; Ye & Xu, 2023; Hazimeh et al., 2021). In addition to the works of Obando Ceron* et al. (2024) and Willi* et al. (2024) already discussed, Farebrother et al. (2024) proposed replacing regression loss with classification loss and showed its benefits on MoEs.
@@
Scaling reinforcement learning (RL) networks presents significant challenges, largely due to issues with training instabilities (Hessel et al., 2018). To address these difficulties, several techniques have been proposed, with a focus on algorithmic improvements or hyperparameter optimization (Farebrother et al., 2022; 2024; Taiga et al., 2022; Schwarzer* et al., 2023). Recently, a few studies have begun to focus on architectural innovations as a means of addressing these challenges. Kumar et al. (2023a) showed that incorporating a learned spatial embedding with the output of a ResNet encoder, prior to flattening it, improves the scalability in offline RL. Obando Ceron et al. (2024) demonstrated that gradual magnitude pruning enables the scaling of value-based agents, resulting in substantial performance gains.
@@
It has been observed that networks trained on non-stationary distributions tend to lose their expressivity over time, a phenomenon referred to as the loss of plasticity (Kumar et al., 2021; Lyle et al., 2021). While several studies have attempted to explain this phenomenon (Lyle et al., 2023; Ma et al., 2023; Lyle et al., 2024), the underlying causes remain uncertain. In a parallel line of research, numerous methods have been proposed to improve network trainability. These approaches involve using small batch size for gradient updates (Ceron et al., 2023), periodically resetting some of the network’s parameters to random values (Sokar et al., 2023; Nikishin et al., 2022; Igl et al., 2020; Dohare et al., 2024), slightly modifying the existing weights by shrinking and perturbing them (D’Oro et al., 2022; Schwarzer* et al., 2023), and restricting the large deviation from initial weights (Kumar et al., 2023b; Lewandowski et al., 2024). Recently, Lee et al. (2024) have also demonstrated that combining reinitialization with a teacher-student framework can enhance plasticity. Another technique involves preserving useful neurons by perturbing gradients to ensure smaller updates for important neurons (Elsayed & Mahmood, 2024). In this work, we explore some of these methods to improve plasticity and utilization of experts.
@@ 8 Conclusions
We began our investigation by evaluating the components that make up a SoftMoE architecture, so as to better understand the strong performance gains reported by Obando Ceron* et al. (2024) when applied to value-based deep RL agents. One of the principal, and rather striking, findings is that tokenizing the encoder outputs is one of the primary drivers of the performance improvements. Since tokenization is mostly necessary for pixel-based environments (where encoders typically consist of convolutional layers), it can explain the absence of performance gains observed by Obando Ceron* et al. (2024) and Willi* et al. (2024) when using MoE architectures in actor-critic algorithms, since these were evaluated in non-pixel-based environments. Initial experiments illustrated in Figure 22 suggest that our findings carry over to actor-critic agents on continuous control environments.
More generally, our findings suggests that the common practice of flattening the outputs of deep RL encoders is sub-optimal, as it is likely not maintaining the spatial structure output by the convolutional encoders. Indeed, the results with PerConv versus Shuffled tokenization in Figure 7 demonstrate that maintaining this spatial structure is important for performance. This is further confirmed by our results in Figure 5, where we can observe that not flattening the convolutional outputs can lead to significant improvements. This idea merits further investigation, as the results vary depending on the architecture used and whether features are summed or averaged over.
@@ Acknowledgements
The authors would like to thank Gheorghe Comanici, Jesse Farebrother, Joan Puigcerver, Doina Precup, and the rest of the Google DeepMind team for valuable feedback on this work. We would also like to thank the Python community (Van Rossum & Drake Jr, 1995; Oliphant, 2007) for developing tools that enabled this work, including NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007) and JAX (Bradbury et al., 2018).
@@ B.1 Comparison between token choice routing and expert choice routing
Figure 12 demonstrates that expert choice routing consistently outperforms token choice routing across varying number of experts. Moreover, this performance gap widens as the number of experts increases, highlighting the advantages of expert choice routing in addressing the limitations of token choice routing.
@@ B.2 Don’t flatten tokenize
In this section, we provide further analysis on the impact of tokenization and their weight combination on models with varying number of experts. To this end, we investigate SoftMoE-2 and SoftMoE-8 and compare it against the scaled SoftMoE-1 and ExpertChoice-1 (non-combined tokenization). Figure 13 shows that single scaled expert can reach a performance close to multiple experts, even in the case of large number of experts. In addition, consistent with our results, combined tokenization adds benefits over the sparse one. Figure 14 shows a comparison between the learning curves of single scaled expert and multiple ones.
Figure 13: Combined tokenization plays a crucial role in performance. SoftMoE-1 with scaled hidden layer achieves a comparable performance to the standard SoftMoE.
@@
To further assess the effect of flattening in reducing performance of agents, we replace the flattening operation in the baseline by global average pooling. We performed these experiments on the baseline with the default dimensionality and the scaled one in Rainbow.
Figure 15: Replacing the flattening operation by global average pooling improves the performance of the baseline and enables scaling.
B.3 Analysis of SoftMoE components
In this section, we present an analysis of each of SoftMoE components on the DER algorithm. As shown in Figure 16, we observe similar tends to our findings in Section 4.2 on Rainbow. Combined tokens offers an additional benefits over the sparse ones. Experts do seem not be specialized in subset of the tokens. Adjusting the architectural depth and width for direct comparison with the baseline doesn’t explain the performance gains of SoftMoE.
Figure 16: Understanding the impact of SoftMoE components on DER.
Effect of Expert width
In section 5, we studied the effect of scaling up the size of the expert hidden layer of SoftMoE-1. Here, we present the results of scaling down the size of this layer by 4. As shown in Figure 17, SoftMoE-1 with scaled down layer still outperform the baseline despite that it has less parameter count.
Figure 17: The effect of scaling down the size of the expert hidden layer in SoftMoE-1.
B.4 More agents
@@ B.5 Expert Utilization
B.5.1 Some experts are redundant
In Section 4.2, we showed that experts are not specialized in learning subset of tokens. We expand this analysis by assessing the redundancy of experts. To this end, we prune two experts of SoftMoE-4 during training. We studied two pruning schemes: pruning the two experts once and gradually prune throughout training. As shown in Figure 19, pruning these experts does not change the performance of the agent.
@@ B.5.1 Some experts are redundant
B.5.2 Improving Expert Utilization
Experimental details
For hyper-parameter tuning, we used six games (Asterix, Demon Attack, Seaquest, Breakout, Beam Rider, Space Invaders). For the reset period, we searched for Rainbow over the grid [$0.5\times 10^{4}$, $10\times 10^{4},50\times 10^{4},100\times 10^{4},125\times 10^{4},500\times 10^{4},1000\times 10^{4}$] gradient steps. The best found is $125\times 10^{4}$, which corresponds to 20M environment steps. For S&P, we searched over the values [(0.5, 0.5), (0.4, 0.1), (0.8, 0.2)] for shrink and perturb hyperparameters respectively, as studied in (Schwarzer* et al., 2023). We used values of (0.5, 0.5).
@@ B.5.2 Improving Expert Utilization
In investigating current techniques to improve expert utilization, we explore reset either experts weights only or both experts and router weights. As illustrated in Figure 20, both methods do not help in improving utilization in Rainbow and DQN, with resetting the router weights leads to worse performance.
Figure 20: Comparison between resetting experts weights only versus experts and router weights in Rainbow (top), DQN (middle), and DER (bottom).
Appendix C Extra environments
To evaluate the generality of our claims, we conducted experiments with Rainbow on Procgen (Cobbe et al., 2019) (Figure 21) and with SAC (Haarnoja et al., 2018) on CALE (Farebrother & Castro, 2024).
Figure 21: Evaluating Rainbow on the 16 easy Procgen (Cobbe et al., 2019) environments for SoftMoE-4 and SoftMoE-1 with the penultimate layer scaled by 4x. Normalization was done using the $R_{min}$ and $R_{max}$ scores from Cobbe et al. (2019). Reporting Median, IQM, Mean, and Optimality Gap (Agarwal et al., 2021), where higher is better for the first three. Consistent with the results in the paper, SoftMoe yields improvements, and the scaled SoftMoE-1 is comparable to SoftMoE-4.
Figure 22: Evaluating tokenization on the SAC encoder (Haarnoja et al., 2018) on the CALE (Farebrother & Castro, 2024), as done in Figure 5. In these experiments, we summed over the first dimension after tokenization. Reporting Median, IQM, Mean, and Optimality Gap (Agarwal et al., 2021), where higher is better for the first three. Consistent with the results in the paper, tokenization yields improvements over the baseline.
C.1 Per-game results
C.1.1 Don’t flatten, tokenize!
Figure 23: Per-game results for tokenized Rainbow-lite with CNN network (corresponding to Figure 5).
Figure 24: Per-game results for tokenized Rainbow-lite with Impala network (corresponding to Figure 5).